OpenStudy (anonymous):

x – y = 7 x + y = 5 What is the x-value of their intersection?

OpenStudy (anonymous):

x – y = 7 =x=y+7

OpenStudy (anonymous):

x + y = 5 =x=5-y

OpenStudy (anonymous):

x=6, y=-1